OpenStudy (anonymous):

What is the product of 2 and 2/3 x 6 and 1/6 in simplest form?

OpenStudy (anonymous):

2 and 2/3 = 8/3 6 and 1/6 = 37/6 8/3 * 37/6 = 296/18 = 148/9